OpenStudy (anonymous):

how do I solve : SEC1035

OpenStudy (anonymous):

sec1035 = \[1/\cos 1035\] Can you solve it from there?

OpenStudy (anonymous):

no can you explain more..? sorry ;P

OpenStudy (phi):

is that 1035 degrees?

OpenStudy (anonymous):

yes

OpenStudy (phi):

Everytime you go 360 degrees you are back where you are started. So you have to divide by 360 and use only the remainder.

OpenStudy (anonymous):

okay so it's Sec315*. what do I do now?

OpenStudy (phi):

I would do cos(315) and then flip it after I get the answer secant(x)= 1/cos(x)

OpenStudy (anonymous):

Secant is the reciprocal of cosine, so you take 1/cos315.

OpenStudy (anonymous):

Cosine is x/r, so the x would be \[\sqrt{2}\], and since you are on a unit circle, the radius (r) would be 1.

OpenStudy (anonymous):

(Because the angle of 315 degrees creates a 45-45-90 triangle.)

OpenStudy (anonymous):

so answer would be root(2)?

OpenStudy (phi):

conquers should memorize cos(45) = sin(45) = sqrt(2)/2 also sin(30)= 0.5 cos(30)= sqrt(3)/2

OpenStudy (anonymous):

Right! Do you understand how that works?

OpenStudy (anonymous):

yes yes ! :D! oh god. sorry for making you elaborate all these, thank you guys!

OpenStudy (anonymous):

I enjoy it! You're very welcome. :]